From GlueXWiki
Jump to: navigation, search

January 2, 2014 FDC meeting


  1. Installation status and plans (Lubomir)
  2. Engineering (Bill)
  3. Electronics (Chris, Nick)
  4. Other


Participants: Eugene, Bill, Nick, Chris, and Lubomir.


- Lubomir on the installation progress since the last meeting: all LV cables are connected to the detector and tested. The test was done by temporary connecting them to the LV panels. Now the cables are grouped in four big bundles (one per quadrant) and stored inside the mesh so that there's space to continue with the signal cables. Some of the labels on the pre-amps had to be corrected today and now we are ready to proceed with the signal cables. We will start with one quadrant of one package, test the connections and then continue with the next package. It is not clear yet if we want to connect the cables to the electronics before proceeding with the next package, but will decide as we go with the cabling.

- The HV was tested during the week just before the holidays: using 90/10 Ar/CO2 pre-mixed bottle and a flow of 400 ccpm as set on the mass flow controller; the packages were supplied in parallel. During the test three CAEN HV modules had to be replaced. The test was done by looking at one long strip per cell with a scope and turning on one HV sector at a time. We saw signals from all the HV sectors. Two of the HV channels in package 1(C1+HV3 and C4+HV2) were tripping every several hours (at +1600V). Another four HV channels (package 2 C6+HV3, C6+HV4 and package 4 C6+HV1, C6+HV2) started tripping when left over the night during the last days of the test. Different reasons for the trips were discussed: HV module, low flow (100 ccpm per package) combined with not so stable gas mixture (most likely), some bad cable connections due to the work on the LV connections and the installation of the prism holders. Plan for further tests: apply now HV since the FDC is being flushed with Nitrogen since the beginning of the Holidays; switch to 90/10 and test these with the Bertan HV supply (for that Nick will make a cable with SHV connector to be connected to the rectangular HV connector). Bill will work on a diving board to be used inside the mesh in case we want to access the individual HV connectors.

- We are also not sure if the mass flow controller works correctly. It is set to 400 ccpm but reads 459 ccpm. We measure the flow also with a rotameter which shows ~0.85 cfph (which is ~400ccpm) when using Ar/CO2 but shows 0.35 cfph when using Nitrogen.


- The cooling system was off during the Holidays. When we turned it on this morning the Fluorinert level was somewhat lower then before the Holidays. Bill is going to mark and check the level to see if there's some leakage.

- No results yet from the last survey done first week of December; Bill will talk to the survey group.


- Nick and Bill discussed the installation of the grounding cable, also in the Hall after the meeting.

- FDC cabling now has first priority, so Nick will make sure that the FEL people are available.